OpenStudy (anonymous):

Two skaters race in different directions. The first skater travels at 18 mph and the second travels at 23 mph. How long will it be before they are 123 miles apart?

jimthompson5910 (jim_thompson5910):

first skater: d=rt d1=18t second skater d=rt d2=23t Add the distances to get d1+d2 =18t+23t =41t So their total distance apart at time t is 41t We want this distance to be 123 miles, so 41t = 123 t = 123/41 t = 3 So it will take 3 hours for them to be 123 miles apart.